OpenStudy (elleblythe):

If csc (x-5pi)=3 then csc x=?? How to get the answer and what is the concept behind this?

OpenStudy (dumbcow):

first recall that csc = 1/sin \[\sin(x-5\pi) = \frac{1}{3}\] next use angle sum identity \[\sin(a-b) = \sin a \cos b - \sin b \cos a\] \[\rightarrow \sin x \cos 5\pi - \sin 5\pi \cos x = \frac{1}{3}\] from unit circle we know that cos(5pi) = cos(pi) = -1 sin(5pi) = sin(pi) = 0 \[\rightarrow \sin x (-1) - 0 = \frac{1}{3}\] \[\sin x = -\frac{1}{3}\] \[\csc x = -3\]